A line has a slope of "m" = (-1/3) and passes through points (-4,k) and (2,3). Find the value of "k"
Answer by nerdybill(7384)

A line has a slope of "m" = (-1/3) and passes through points (-4,k) and (2,3). Find the value of "k"
.
Slope between any two points is:
m = (y2 - y1)/(x2 - x1)
.
plugging in what was given in the problem:
-1/3 = (3 - k)/(2 - (-4))
-1/3 = (3 - k)/(2 + 4)
-1/3 = (3 - k)/6
Cross-multiplying:
3(3-k) = -6
3-k = -2
-k = -5
k = 5